Fermi National Laboratory
 

Operation Summary for October 11, 2010 - October 13, 2010

 

Monday Oct. 11
The day shift began with Operations monitoring the Tevatron (TeV) store 8162 (330.76E30) and the Recycler stash (8.7E10), with the Antiproton Source (Pbar) stacking (62.1E10), with the Linac, Booster, and the Main Injector (MI) operating normally, with NuMI off for target work, with the MiniBooNE experiment taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, operators monitored the machines.

During the evening shift, Operations terminated store 8162 at 9:06 PM. Operators reported a TECAR problem, which required some recovery. Operations put the TeV into HEP shot setup at 11:33 PM, but the setup was put on hold due to MI RF problems.

Tuesday Oct. 12
The midnight shift began with MI RF problems, with Operations monitoring the stash (353.1E10) and the Pbar stack (79.6E10), and with MiniBooNE taking beam.

During the midnight shift, RF personnel resolved the MI RF problems at 1:21 AM. Operations put the TeV into shot setup, but at 2:30 AM the setup ended in a quench during the ramp. Operators prepared D0 for access, and then began the TeVcryo system recovery, turn-on, and checkout.

During the day shift, operators prepared MiniBooNE for access at 9:34 AM. Operations completed the TeV checkout at 11:10 AM, put the TeV into shot setup, and established store 8170 at 12:27 PM with an AIL of 269.29E30. MiniBooNE came out of access at 1:34 PM.

During the evening shift, MRF9 failed at 6:13 PM.

Wednesday Oct. 13
The midnight shift began with Operations monitoring store 8170 (56.05E30) and the stash (314.0E10), with Pbar stacking (10.4E10), and with MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8170 at 2:08 AM, put the TeV into shot setup, and established store 8171 at 3:36 AM with an AIL of 297E30. Operations did not transfer antiprotons to the Recycler following the store. Operators prepared MI for access at 7:26 AM to repair MIRF9 and MIRF15.

The Plans for Wednesday
The plans for today are to monitor the store, while the rest of the complex conducts maintenance during the MI RF repairs.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/13/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory